OpenStudy (anonymous):

Calculate the second derivative of the function. f(θ) = sec^2(2θ)

OpenStudy (anonymous):

first derivative is 4tan(2theta)sec^2(2theta) right?

OpenStudy (helder_edwin):

yes. it is right.

OpenStudy (helder_edwin):

so u now have \[ \large f'(\theta)=4\sec^2(2\theta)\tan(2\theta) \] differentiate again.

OpenStudy (anonymous):

this is what I got

OpenStudy (helder_edwin):

it is wrong

OpenStudy (helder_edwin):

i got \[ \large f''(\theta)=16\sec^2(2\theta)\tan^2(2\theta)+8\sec^4(2\theta) \]

OpenStudy (anonymous):

= 16 sec² (2x) tan² ( 2x) + 8 sec^4 (2x)

OpenStudy (anonymous):

It'll be easier if you leave out the constant before taking derivative: = 4 [ 4 sec² (2x) tan² ( 2x) + 2 sec^4 (2x) ]

OpenStudy (anonymous):

Oh, yeah that's helpful. I understand :) thank you

OpenStudy (helder_edwin):

don't ever forget this rule: \[ \large (\alpha f(x))'=\alpha f'(x) \] usually, applying it easies things.

OpenStudy (anonymous):

I don't understand that notation though, the other comment helped a lot

OpenStudy (anonymous):

\=? large=?